A sound wave has a frequency of 2500 Hz and a speed of 1500 m s–1. 
 What is the shortest distance from a point of maximum pressure in the wave to a point of minimum pressure?

Grade:12

1 Answers

Shrey Thakkar
6 years ago
Frequency=2500 hz
Wave speed=1500m/s
Now wave speed = wave length*Frequency
So wavelength=Wave speed/Frequency
Wavelength=1500/2500
                   =3/5
                   =0.6 m
Sound wave is also known as a pressure wave. The wavelength can be said to be the distance between 2 consecutive maximum pressure points or distance between 2 consecutive minimum pressure points.
In the middle of 2 consecutive maximum pressure points, there is a minimum pressure point.
So wavelength between point of maximum pressure in the wave to a point of minimum pressure is=0.6/2
                                                                                                                                                            =0.3 m
Hence the shortest distance between point of maximum pressure in the wave to a point of minimum pressure is 0.3m.